OpenStudy (anonymous):

Solve the system by substitution: 2x+3y=20 6x-y=20

OpenStudy (anonymous):

2x + 3y=20 2x=20-3y ( BY TRANSPOSING 3y to RHS) 2x/2=(20-3y)/2 x=(20-3y)/2 PUT THE VALUE OF x IN THE EQUATION 2 : 6x-y=20 6((20-3y)/2) - y=20 (120-18y)/2 - y=20 (120-18y - 2y)/2=20 120-20y=40 transposing 120 to RHS -20y=-80 y=4 and 6x-4=20 6x=24 x=4 hence x and y are equal to 4

OpenStudy (anonymous):

Thank you! I have another one to ask